5x=3(1-y)
3x+3y+1=0

Answer by fractalier(6550)

From
5x=3(1-y)
3x+3y+1=0
let's multiply the first one out and get
5x = 3 - 3y
If we solve it for 3y, we can plug that into the other equation...
3y = -5x + 3
and
then
3x + (-5x + 3) + 1 = 0
-2x + 4 = 0
-2x = -4
x = 2
Then substitute and get
3(2) + 3y + 1 = 0
3y + 7 = 0
3y = -7
y = -7/3